### v1

# Temporally Consistent Factuality Probing for Large Language Models

Abstract
The prolific use of Large Language Models (LLMs) as an alternate knowledge base requires them to be factually consistent, necessitating both correctness and consistency traits for paraphrased queries. Recently, significant attempts have been made to benchmark datasets and metrics to evaluate LLMs for these traits. However, structural simplicity (subject-relation-object) and contemporary association in their query formulation limit the broader definition of factuality and consistency. In this study, we introduce TeCFaP, a novel Temporally Consistent Factuality Probe task to expand the consistent factuality probe in the temporal dimension. To this end, we propose TEMP-COFAC, a high-quality dataset of prefix-style English query paraphrases. Subsequently, we extend the definitions of existing metrics to represent consistent factuality across temporal dimension. We experiment with a diverse set of LLMs and find most of them performing poorly on TeCFaP. Next, we propose a novel solution CoTSeLF (Consistent-Time-Sensitive Learning Framework) combining multi-task instruction tuning (MT-IT) with consistent-time-sensitive reinforcement learning (CTSRL) to improve temporally consistent factuality in LLMs. Our experiments demonstrate the efficacy of CoTSeLF over several baselines.
**footnotetext: Equal contribution
1 Introduction
The emergence of large language models (LLMs) significantly contributes to achieving broader objectives of Artificial General Intelligence (AGI) in acquiring self-learning capabilities to solve complex tasks Ge et al. (2023). As the landscape of LLMs is continuously evolving, the capability to comprehend, process, and reason about time, known as temporal reasoning, has emerged as one of the cornerstones for advanced cognitive processing in LLMs. Temporal reasoning goes beyond greater recognition of dates and duration; it embodies the convoluted understanding of temporal dependencies, chronological sequences, and the nuanced, consistent relationship of events across time.
Why is consistent factuality important? Knowledge bases (KBs) were the foremost choice in factual knowledge retrieval tasks before the appearance of LLMs. A KB is a structured database containing a collection of facts (subject, relation, object) Lan et al. (2021). There has been a surge of interest in using pre-trained LLMs as KBs. AlKhamissi et al. (2022) presented an extensive review on significant developments Petroni et al. (2019) Dhingra et al. (2022) Heinzerling and Inui (2021) in this direction. One of the biggest appeals of using LLMs as KBs is that a query can be written in natural language instead of relying on a specific KB schema Elazar et al. (2021). Due to the complex nature of language, the semantic meaning can be expressed in multiple surface forms. Accurate and consistent retrievals, despite a change in surface form, are two fundamental traits required from LLMs to replace KBs. Built on a manually engineered schema that dictates the possible set of entities and relations, KBs ensure factual and consistent answers AlKhamissi et al. (2022). On the contrary, inconsistent factuality is reported as a widespread problem in LLMs, especially in autoregressive setting Tam et al. (2022).
Probing consistent factuality. Usually, factuality (accuracy) is used as a widespread metric in probing LLMs to check linguistic capabilities such as commonsense Zhang et al. (2020); Forbes et al. (2019) and reasoning Talmor et al. (2020); Kassner et al. (2020). On the other hand, an increase in public access of LLMs requires them to consistently respond to user-specific diversities in surface forms of a query. Ravichander et al. (2020) measured consistency through paired probes. Elazar et al. (2021) extended the work by investigating and improving the consistency of LLMs behavior across different factual knowledge types.
Our novel task and dataset. So far, the consistent factuality probing in LLMs has been predominantly contemporary – the existence of a subject and an object associated via a relation is coeval, i.e., a typical query from PARAREL dataset Elazar et al. (2021) is – X was born in Y in which subject X (person) and object Y (location) are connected via relation born-in. Note that the subject and object are contemporaneous; both exist simultaneously. As information keeps on getting generated, maintained and lost over time, the above formulation is insufficient in capturing the temporal association between the queried entity (subject) and expected value entity (object), therefore providing a poor representation of the overall consistent factuality of LLMs.
To address this, we present Temporally Consistent Factuality Probe (TeCFaP), a novel task accompanied by a new dataset, TEMP-COFAC. TeCFaP seeks to exploit the temporal association among entities via a subject-relation pair to represent temporally consistent factuality of LLMs. It defines the query structure in the form of (key_object, subject-relation, value_object). The proposed formulation expands the probe in the temporal dimension. As we observe in Figure 1, the space is three-dimensional – subject, relation, and time. Since time has a directional attribute, we expect temporal association to be either in a forward or backward direction. For example, the query Hybrid Theory was released by linkin park just before [Meteora] is a forward direction probe where a key_object (Hybrid Theory) placed at time t is temporally associated with a value_object (Meteora) in time-space at $(t+1)$ via a subject (Linkin Park) and a relation (release-by). At present, our probe is limited to only strict associations where the expected value_object is either located at $(t+1)$ or $(t-1)$ step in temporal space w.r.t key_object at time $t$. Inspired by Pustejovsky et al. (2005), strict association is achieved via trigger words such as immediately before, right after, soon after, etc. We observe the destitute performance of LLMs on TeCFaP metrics (defined in Section 3) – temporal factuality, temporal consistency, and temporally consistent factuality are in [$0.95\%$ $3.63\%$], [$13.28\%$ $64.87\%$] and [$0\%$ to $2\%$] range, respectively in zero-shot setting.
Figure 1: Symbolic representation of the TeCFaP  objective. An entity key_object holds a temporal relationship with another entity value_object via a subject-relation pair in either direction – forward or backward.
Our proposed model. We present CoTSeLF, a Consistent-Time-Sensitive Learning Framework, built on a multi-task instruction-tuned (MT-IT) framework followed by consistent-time-sensitive reinforcement learning (CTSRL) to improve temporally consistent factuality in LLMs. We compare CoTSeLF with several baselines and show it outperforming the best baseline Tan et al. (2023) by $12.7\%$, $10.9\%$ and $90.4\%$, respectively, for temporal factuality, temporal consistency, and temporally consistent factuality.
Contributions. In short, we make the following contributions through this study***We will opensource the source code and dataset.:
  • We establish the need for temporally consistent factuality in LLMs and propose TeCFaP, a novel task (Section 3).
  • We create TEMP-COFAC, a novel dataset consisting of $66$ diverse subject-relation pairs and $8$ paraphrase samples each for forward and backward temporal association for a given subject-relation pair (Section 2).
  • We experiment with a diverse set of LLMs. Our experiments and analyses highlight how LLMs poorly perform on TeCFaP (Section 5).
  • We propose CoTSeLF, a framework to improve temporally consistent factuality in LLMs (Section 4). Our experiments highlight that CoTSeLF surpasses the recent baseline models (Section 5). We further analyze how the probabilistic space evolves under CoTSeLF (Appendix A.2.4).
Figure 2: The architectural framework of TEMP-COFAC – (1) a set of diverse subject-relation pairs, (2) a sequence of entities which are temporally connected via a given subject-relation pair, (3) a set of paraphrase templates with a placeholder for key_object and value_object developed from subject-relation pairs, and (4) a closed vocabulary candidate set developed from possible entity space for a given subject-relation pair.
2 The TEMP-COFAC Dataset
Here, we present a novel English prefix-style TEMP-COFAC dataset with a temporal range of $1526$-$2022$. Inspired by Elsahar et al. (2018), we semi-automatically***We leverage GPT-4 to extend manually constructed initial set of base subject-relation pairs. curate a diverse set of base subject and relations pairs subject-relation. Next, we define $E_{i}$, a strict temporally ordered set of entities associated with $i^{th}$ subject-relation pair. An entity can act as a key_object or a value_object relative to the role of another entity positioned right next/before it. We then create base_patterns for both forward and backward directions following Petroni et al. (2019). Afterwards, a set of patterns $P_{i}$ is constructed by employing paraphrasing techniques Bhagat and Hovy (2013) on base_patterns in forward and backward directions followed by a candidate set $C_{i}$.
Construction approach. The TEMP-COFAC resource is constructed by three NLP experts***All of them are male with their ages ranging between 25-35 years. with a mean cross-annotator agreement of $4.84\pm 0.39$ and $4.86\pm 0.49$ out of $5$ maximum on Likert scale Likert (1932) for factuality and consistency, respectively (refer Appendix A.1 for more details). Following Elazar et al. (2021), our construction process broadly follows a four-step procedure described in Figure 2.
(i) First, we define a set of $m$ diverse subject-relation pairs randomly collected from varied domains – entertainment, technology, politics, automobiles and corporate. Using annotators’ linguistic expertise, we then define base_patterns for forward and backward directions, i.e., Linkin Park released [X] just before [Y] and, Linkin Park released [X] just after [Y] are two base_patterns examples of the forward and backward temporal associations, respectively, where [X] and [Y] are the placeholders for a key_object and a value_object, respectively.
(ii) Next, we manually curate a set of entities $E_{i}$ ($\forall i=0,\ldots,m$) for the $i^{th}$ subject-relation pair such that they have temporal association with it in ascending temporal sequence $t=0,\ldots,j$, where, entity $e_{i}^{t}$ precedes $e_{i}^{t-1}$ and is followed by $e_{i}^{t+1}$ in the temporal space. The cardinality of $E_{i}$, represented by $j$, varies for each subject-relation pair.
(iii) A set of paraphrased patterns $P_{i}$ ($\forall i=0,1,\ldots,m$) is constructed through an application of an online paraphraser tool, Quillbot***https://quillbot.com/ on base_patterns. Set $P_{i}$ defines $r$ and $s$ number of paraphrases in the forward and backward directions, respectively. For simplicity, here we consider uniform values of $r$ and $s$ to be $8$.
(iv) Finally, a constrained candidate set $C_{i}$ is developed as an unordered set of words from $E_{i}$.
Table 1 summarizes the statistics of TEMP-COFAC. Readers can refer to Appendix A.1 for more detail about TEMP-COFAC, including annotation quality, temporal and entity type distributions.
3 TeCFaP Task Structure
We pose TeCFaP  as a sentence completion task that aligns with the general objective of an LM. With prefix style, paraphrase templates for a subject_relation pair are filled with only key_object, and we expect the model to generate value_object. Next, we define TeCFaP metrics to evaluate LLMs.
Temporal factuality and temporal consistency. We extend the metrics defined by Elazar et al. (2021) to TeCFaP. The first metric, temporal-factuality, captures the accuracy of a model across the temporal direction. In contrast with their definition of exact-match accuracy, we define factuality as soft accuracy, a ratio of the number of continuous matches of words in actual and generated value_object for a sample. It helps capture the partially correct generation as well. Next, we define its sub-classification across temporal directions. The forward temporal-factuality measures the accuracy in the forward temporal direction where value_object is located at $t+1$ time step for a given key_object at $t$. In backward temporal-factuality, the value_object is located at $t-1$ time step for a given key_object at $t$.
# Subject-Relation pairs 66
# Paraphrase patterns 1056
        # Forward patterns 528
        # Backward patterns 528
Avg # pattern per relation-subject 18
# Entities 700
        #Unique entity types 11
        Min # entities per relation-subject 2
        Max # Entities per relation-subject 16
        Avg # entities per relation-subject 10.6
# Unique samples in dataset 10144
Table 1: High-level statistics of the TEMP-COFAC dataset.
The second metric is temporal-consistency. Given a pair of prefix-style paraphrases for a subject_relation filled with an identical key_object, an identical value_object should be generated. The metric estimate is binary (one or zero) for a given pair of such paraphrases if the model’s responses are identical or different. Forward and backward directions paraphrases contribute to forward and backward temporal-consitency, respectively.
Temporally consistent factuality. The third composite metric, temporally-consistent-factuality, is a stricter version of temporal-factuality, requiring a model to be consistent and factual across the temporal direction. It reports the temporal-factuality only if the responses are identical from all prefix-style paraphrases for a given subject_relation and key_object in a particular temporal direction. Paraphrases in forward and backward directions contribute to forward and backward temporally-consitent-factuality, respectively.
Other metrics. The quality of patterns is measured through temporal-succ_patt, which reports the percentage of patterns that return a correct value_object at least once in the entire probe. Also, to estimate the degree of temporal knowledge in models, we define temporal-succ_objs. Here, we report the percentages of value_objects, which are correctly generated at least once by the model. Further, we define temporal-know_cons and temporal-unk_cons as metrics to measure temporal-consistency of the fraction of patterns which generated correct value_object at least once and the fraction of patterns which never responded with a correct value_object for a subject_relation, respectively. All four metrics are then classified into forward and backward temporal directions.
4 Consistent-Time-Sensitive Learning Framework (CoTSeLF)
Equipped with the advancements in model fine-tuning combined with the recent baseline in temporal reasoning, We begin with a base pre-trained LLM and apply supervised multi-task instruction-tuning to develop a multi-task instruction-tuned (MT-IT) model. Subsequently, we apply time and consistency-sensitive reinforcement learning to the MT-IT model to enhance its temporally consistent factuality capabilities.
Motivation. Instruction tuning (IT) Zhang et al. (2023) is widely used as a low-cost, efficient modeling technique in developing domain or task-specific models. In recent IT advancements, a parameter-efficient fine-tuning-based (PEFT) Mangrulkar et al. (2022) method such as low-rank adaptation (LoRA) Hu et al. (2021) has shown tremendous potential to apply effective instruction tuning on LLMs for low-cost infrastructure settings. Moreover, RL combined with supervised fine-tuning has proven effective in improving the fine-tuned model in several domain-specific tasks, including temporal reasoning Tan et al. (2023). Furthermore, advances in multi-task learning led to several breakthroughs in modeling multiple objectives simultaneously Zhang and Yang (2022).
Multi-Task Instruction-Tuning (MT-IT). We consider a multi-objective optimization problem that simultaneously improves model’s factuality and consistency, denoted by tasks k1 and k2, respectively. In k1, we apply a standard sentence completion task where an incomplete sentence filled with key_object, augmented with a context and task instruction, is passed as input to allow the model complete the sentence with expected value_object. Whereas k2 is a binary task predicting true or false if two sentences are paraphrased. We start with transforming the TEMP-COFAC dataset into an instruction-based dataset in line with the formats proposed by Wang et al. (2023); Taori et al. (2023) (Figure 3). Next, we consider a base pre-trained LLM with parameter $\theta$ and then apply LoRA instruction-based supervised fine-tuning where the objective is to maximize ${p(o^{k1}|s^{k1},i^{k1},c)}$ for improving model’s factual capability, where $s^{k1}$, $i^{k1}$, $c$ and, $o^{k1}$ are instruction, input, context and output, respectively for task k1. We further add another objective of maximizing ${p(o^{k2}|s^{k2},i^{k2})}$ to improve the model’s consistency in extending the framework to multi-task learning setup (Figure 3), where $i^{k2}$ is a pair of sentences randomly sampled positively and negatively to predict the output $o^{k2}$ as true or false, respectively.
Figure 3: An instruction-based sample from training data for MT-IT model. Task k1: Generative sentence completion; Task k2: Binary paraphrase prediction.
Consistent Time-Sensitive Reinforcement Learning (CTSRL). Multiple value_objects for a given key_object are possible by excluding the temporal direction. We propose CTSRL to model consistent sensitivity towards time and overcoming the limitation of binary temporal character defined by TSRL Tan et al. (2023). To further fine-tune $\theta$, we apply CTSRL to model consistent temporal sensitivity by jointly modeling k1 and k2. The reward function is the linear combination of temporal and consistent sensitivities components. First, we define $\texttt{CTSRL}_{Discrete}$ in line with the original reward model of TSRL, where a positive discrete reward of value equal to one is given for correct prediction and zero for a negative answer. We also define $\alpha$ as the weight parameter for the consistence sensitivity reward component.
Models Temp-fact Temp-cons Temp-cons-fact
Avg Bwd Fwd Avg Bwd Fwd Avg Bwd Fwd
GPT-J [6B] 3.63 6.88 0.37 64.87 66.29 63.46 1.48 2.73 0.22
Falcon [7B] 2.99 5.74 0.24 41.52 40.42 42.63 1.08 2.03 0.13
LLaMA [7B] 1.48 0.89 2.07 13.28 12.52 14.03 0 0 0
LLaMA [13B] 1.11 1.01 1.21 15.65 16.33 14.96 0.2 0.17 0.24
LLaMA2 [7B] 0.95 0.73 1.17 22.34 21.95 22.73 0.08 0 0.17
LLaMA2 [13B] 1.13 0.97 1.27 13.34 14.41 12.26 0.09 0 0.19
Table 2: Zero-shot performance in open vocabulary setting on TeCFaP across various LLMs. Temp-fact: temporal factuality (in %), Temp-cons: temporal consistency (in %), Temp-cons-fact: temporally consistent factuality (in %). Fwd and Bwd are the forward and backward direction probe, respectively (Avg, an average of both directions).
$R_{d}(x)=(1-\alpha)R_{d}^{t}(x)+\alpha R_{d}^{c}(x)$ (1)
$R^{\lambda}_{d}(x)=\begin{dcases}P^{\lambda}_{d}(x),&\textit{if }{O_{g}}(\theta(x))=O_{l}(x),\\ N^{\lambda}_{d}(x),&otherwise.\end{dcases}$ (2)
For a given input $x$, the overall reward function $R_{d}(x)$ in the $\texttt{CTSRL}_{Discrete}$ setting is presented in Equation 1, where $R_{d}^{t}(x)$ and $R_{d}^{c}(x)$ are reward contribution for tasks k1 and k2, respectively. In Equation 2, $\lambda$ is used as a task indicator where $P^{\lambda}_{d}(x)$ is a positive reward score for consistence sensitivity component if $\lambda$ equals to $c$. Similarly, $N^{\lambda}_{d}(x)$ is a negative reward score. ${O_{g}}(\theta(x))$ and $O_{l}(x)$ are the generated value_object and ground-truth label value_object, respectively for given input $x$. In the case of $\texttt{CTSRL}_{Discrete}$, a positive reward score equal to one is assigned in case of correctly generated output and zero otherwise for both tasks k1 and k2. We further define another variant, $\texttt{CTSRL}_{Smooth}$, to model continuous and relative properties of time. The temporal sensitivity reward is a continuous function where a positive reward has a maximum function value equal to one. However, the negative reward is proportional to the relative distance of the predicted answer from the correct answer in the temporal axis. The goal is to penalize incorrect answers that are distant from the correct answer more severely than the incorrect predictions that are nearby on the temporal axis. There is no change in the reward component for consistence sensitivity except for releasing the constraint of the parameter $alpha$, indicating that $\texttt{CTSRL}_{Smooth}$ is an unweighted linear combination of continuous temporal and discrete consistence sensitivity reward components.
$R_{s}(x)=R_{s}^{t}(x)+R_{s}^{c}(x)$ (3)
$N^{t}_{s}(x)=\begin{dcases}\frac{|t_{O_{l}}-t_{O_{g}}|}{t_{n}-t_{O_{l}}},&\textit{if }t_{O_{g}}>t_{O_{l}},\\ \frac{|t_{O_{l}}-t_{O_{g}}|}{t_{O_{l}}},&\text{otherwise}\end{dcases}$ (4)
Equation 3 presents reward function $R_{s}(x)$ in $\texttt{CTSRL}_{Smooth}$ for a given input $x$, where $R_{s}^{t}(x)$ and $R_{s}^{c}(x)$ are the reward contribution for tasks k1 and k2, respectively. For the temporal sensitivity task (k1), the positive reward score is assigned as a value equal to one in case of correctly generated output, and the negative reward score is the relative distance of the wrong prediction from the correct prediction in the temporal axis as presented in Equation 4. The symbol $t_{O_{l}}$ is the time step of the ground-truth label, $t_{O_{g}}$ represents the time step for generated output, and $t_{n}$ is the end time step of entity sequence for that subject-relation pair.
Figure 4: Results for temporally consistent factuality (Temp-cons-fact) in $k$-shot ($k$=1,2,3) ICL setup with LLaMA[13B] in an open vocabulary setting.
5 Experimental Results
Experimental Setup. The families of GPT-J***https://huggingface.co/EleutherAI/gpt-j-6b, Falcon Almazrouei et al. (2023), and LLaMA Touvron et al. (2023) are considered for evaluation on TeCFaP. Primarily, we evaluate LLMs in an open vocabulary setting where the next token is sampled from the entire vocabulary. Next, we conduct experiments in an in-context learning setup, followed by a closed vocabulary setting. Finally, the CoTSeLF efficacy evaluation is conducted.
Temporally consistent factuality on TeCFaP. We start with a comparison of various LLMs in zero-shot setting. The task is to correctly complete a sentence with the expected value_object given an instruction followed by an input, i.e., "complete the given sentence with the correct phrase: Meteora was released by Linkin Park immediately after". We observe the destitute performance of all the experimented LLMs over both temporal-factuality and temporally-consitent-factuality in the range of [$0.95\%-3.63\%$] and [$0\%-1.48\%$], respectively (Table 2). We notice that GPT-J and Falcon tend to be highly consistent in the range of [$41.52\%-64.87\%$] while being miserably factually incorrect compared to the LLaMA model in the same range of parameters size. Additionally, we notice that various families of LLMs behave differently regarding their sensitivity towards temporal direction. However, temporal direction and temporal consistency do not significantly correlate. Further, a preliminary evaluation of TeCFaP in zero-shot setting for GPT-4 OpenAI et al. (2023) with a minimal sample size yields temporal-factuality of $31.0\%$. (Appendix A.2.6).
Figure 5: Average temporally consistent factuality and temporal factuality (second y-axis) in an open vocabulary and two-shot setting across temporal bins of Entities (bin size: $10$ years) with LLaMA[13B].
In-context setup. In-context Learning (ICL) helps off-the-shelf LLMs solve unseen tasks without the requirement of fine-tuning Dong et al. (2023). We provide $k$ randomly-drawn examples from the same subject_relation pair as supplementary context in a $k$-shot ICL setting, i.e., a one-shot example is as follows: "complete the given sentence with the correct phrase: Meteora was released by Linkin Park immediately after => Hybrid Theory. American band LP released Minutes to Midnight immediately after =>". This test evaluates LLaMA[13B] and varies $k$ in the range [1-3]. In a two-shot setup, we observe absolute percentage points improvement of $1.76$ in temporally-consistent-factuality (Figure 4). At the same time, the improvements of $7.01$ and $18.22$ percentage points are noted in contributory metrics temporal-factuality and temporal-consistency, respectively (refer to Appendix A.2.1 for more details).
Figure 5 presents temporal-factuality across the temporal distribution of entities. Findings reveal that temporal-factuality for entities belonging to the historical period ($1500$-$1800$) is significantly higher, with an average of $25.27\%$ compared to $9.08\%$ for the entities in the contemporary period (after $1800$). At the same time, the presence of multiple sources of the same information in LLMs pre-trained dataset for the contemporary period leads to better temporally-consistent-factuality.
Figure 6: Qualitative results in an open vocabulary with two-shot ICL setup across LLaMA variants. Error bars represent divergence across temporal directions – forward and backward.
Additionally, in Figure 6, qualitative analysis reveals that LLaMA[7B] attains best temporal-know_cons at $43.29\%$ with a divergence of $10.00$ percentage points between known and unknown temporal consistencies across LLaMA variants. On the other hand, $35.04\%$ patterns yield a correct value_object at least once in contrast to only $16.18\%$ value_objects, which were predicted correctly once in the entire probe for a model.
Closed vocabulary setup. The next word generated by an LM can still be the right placement given its general objective to maximize the semantic expectation irrespective of the expected value_object. Therefore, we also conduct experiments in a closed vocabulary setting where the sample space is reduced to a candidate set***A set of restricted tokens are generated by employing byte pair encoding Sennrich et al. (2016) on candidate set. (defined in Section 2) during generation. This approximation helps set up the probe as KB fact extraction from a given possible facts space, thus maximizing the behavioral expectation of LM as KBs. With an improvement in the range [$1.77\%$ $2.08\%$], LLaMA[13B] has best scores of $1.97\%$ and $3.51\%$ for temporally-consistent-factuality in one and two shots setting, respectively, across variants of LLaMA under closed vocabulary setting (Figure 7). We observe a similar trend for temporal-factuality and temporal-consistency, presented in Appendix A.2.2.
Figure 7: A comparison of average temporally-conistent-factuality between open and closed vocabulary settings across LLaMA variants in zero-shot and 2-shot.
Improvements with CoTSeLF. We conduct all experiments in an open vocabulary setting (assuming no access to candidate sets during inferences) with LLaMA[13B]. First, TEMP-COFAC is vertically split (test ratio: $0.3$) to produce a train set and a test set containing 46 and 26 subject-relation pairs, respectively. The random vertical split ensures the stricter evaluation of CoTSeLF as the test set contains only unseen subject-relation pairs. We broadly categorize this evaluation into four categories: (i) the default performance of the model in zero-shot and two-shot (ICL) setup, (ii) variants of instruction-tuned (IT) models based on the presence of a context along with the novel multi-task IT model, (iii) IT (with context) followed by TSRL; a strong baseline model, and (iv) CoTSeLF, a combined strategy of MT-IT followed by variants of novel CTSRL method. Empirically, we set $\alpha$ as $0.66$ in formulating a discrete variant of CTSRL (significance of $\alpha$ is presented in Appendix A.2.5).
Models Setting (open vocab) Temp-fact Temp-cons Temp-cons-fact
Avg Bwd Fwd Avg Bwd Fwd Avg Bwd Fwd
Default Zero-shot ${0.40}_{\pm 0.00}$ 0.28 0.52 ${16.31}_{\pm 0.00}$ 15.22 17.40 ${0.00}_{\pm 0.00}$ 0.00 0.00
Two-shot [ICL] ${4.95}_{\pm 0.40}$ 4.26 5.63 ${31.57}_{\pm 1.50}$ 29.23 33.91 ${0.53}_{\pm 0.18}$ 0.50 0.57
IT Without context ${11.53}_{\pm 0.08}$ 10.36 12.69 ${25.08}_{\pm 0.06}$ 26.76 23.41 ${2.59}_{\pm 0.30}$ 2.67 2.50
+ Context ${16.87}_{\pm 0.17}$ 16.79 16.96 ${34.39}_{\pm 0.92}$ 34.45 34.31 ${1.75}_{\pm 0.15}$ 2.11 1.38
MT-IT + Context ${18.17}_{\pm 0.15}$ 17.83 18.52 ${36.60}_{\pm 0.36}$ 36.49 36.70 ${3.33}_{\pm 0.41}$ 2.82 3.85
$\Delta$a $1.3\uparrow$ $2.21\uparrow$ $1.58\uparrow$
Baseline IT (+context) + TSRL ${16.60}_{\pm 0.51}$ 17.04 16.15 ${33.23}_{\pm 0.38}$ 33.28 33.18 ${2.28}_{\pm 0.44}$ 2.88 1.67
CoTSeLF MT-IT+$\texttt{CTSRL}_{Discrete}$ $\textbf{18.72}^{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P_{1}}}_{\pm 0.25}$ 19.16 18.29 $\textbf{36.88}^{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P_{2}}}_{\pm 0.49}$ 37.58 36.17 $\textbf{4.34}^{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P_{3}}}_{\pm 0.27}$ 4.45 4.22
MT-IT+$\texttt{CTSRL}_{Smooth}$ ${18.16}_{\pm 0.05}$ 18.11 18.21 ${36.07}_{\pm 0.48}$ 37.09 35.04 ${3.89}_{\pm 0.92}$ 3.74 4.05
$\Delta$b $2.12\uparrow$ $3.65\uparrow$ $2.06\uparrow$
${\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P_{1}=0.003,P_{2}=0.003,P_{3}=0.003}$
Table 3: Experimental results of CoTSeLF across – temporal-factuality, temporal-consistency, and temporally-consistent-factuality (in %), in comparison to multiple baselines on test data with LLaMA[13B] (average scores over three runs). $\Delta$a: improvements of MT-IT over an IT model, $\Delta$b: improvements of CoTSeLF over a baseline model. ($P_{1},P_{2},P_{3}$): $p$-values at CoTSeLF’s best scores compared to baseline model with $n=3$ and one-tailed test.
In Table 3, the additional context provided with an input improves temporal-factuality by $5.22$ percentage points for an IT model. The MT-IT model improves temporal-factuality, temporal-consistency and temporally-consitent-factuality by $7.7\%$, $6.4\%$ and $90.2\%$, respectively, compared to the IT model. We further observe improvements of $12.7\%$, $10.9\%$ and $90.4\%$ in temporal-factuality, temporal-consistency and temporally-consitent-factuality, respectively, with CoTSeLF over the baseline, indicating that improving a model’s temporal consistency also positively impacts its temporal factuality. However, the inaccessibility of GPT-4 architecture limits us to assess the efficacy of CoTSeLF’s on this model. (the probabilistic space evolution under CoTSeLF is presented in Appendix A.2.4).
6 Error Analysis
A causal analysis is conducted to determine any correlation between data characteristics and failure cases. Figure 8 shows a strong correlation between entity-type and temporally-consistent-factuality. Entity types not exclusively attached to a subject-relation pair, such as movie names and geographical locations, perform poorly compared to other entities, such as satellite, person, and software names. It would be interesting to enhance both TEMP-COFAC and CTSRL formulation to include such data characteristics explicitly. For more detail on this, readers can refer to Appendix A.2.3.
Figure 8: Temporally consistent factuality across various entity-types present in test data for $\texttt{CTSRL}_{Discrete}$.
7 Related Work
Consistency in knowledge bases was extensively studied, with developments around data and methods to benchmark the degree of inconsistencies Hansen and Jaumard (2000); Andersen and Pretolani (2001); Thimm (2013) across several tasks; QA Kassner et al. (2020), reading comprehension Antol et al. (2015); Rajpurkar et al. (2016); Ribeiro et al. (2019), summarizing Xie et al. (2021); Roit et al. (2023) and NLI Li et al. (2019).
Notable works in temporal information extraction include TimeBank Pustejovsky et al. (2003) and TimeEval Verhagen et al. (2010). The annotations defined in these datasets are primarily for time-event and events relationships such as before/after. The remarkable advancement in studying temporal reasoning over KGs led to the development of datasets such as TEQUILA Jia et al. (2018a), TimeQuestions Jia et al. (2021), and CronQuesions Saxena et al. (2021), probing KG’s response in ranking entities for a given temporal query. Subsequently, datasets like TEMPLAMA Dhingra et al. (2022) and StreamingQA Liška et al. (2022) were curated for temporal reasoning in LMs. The TSQA Chen et al. (2021) dataset has addressed the drawback of the limited time span of queries, but it defines only time-event relationships. TEMP-REASON Tan et al. (2023) captures multiple facets of temporal reasoning – time-time, time-event, and event-event relationships over a longer time span. However, it does not help evaluate consistent temporal reasoning that our dataset, TEMP-COFAC, addresses. TEMP-COFAC also overcomes a significant limitation of TEMP-REASON’s monotonous anchor queries (anchored only to one entity type; person name as subjects or objects) by providing a diverse set of anchor queries via sixty-six unique sub-rel-obj triplets formed around eleven entity-types. Additionally, it has a broader representation of eleven entity types compared to six in TEMP-REASON (Appendix A.1).
8 Conclusion
This paper presented a new task TeCFaP with a novel resource, TEMP-COFAC, to evaluate temporally consistent factuality in large language models. The contribution continued to present a novel solution CoTSeLF based on multi-task instruction tuning (MT-IT) combined with consistent-time-sensitive reinforcement learning (CTSRL) to improve LLMs temporally consistent factuality. We observed that CoTSeLF outperforms the baselines to improve temporally-consistent-factuality in LLMs. The contribution and findings in this paper would help better understand and enhance LLMs’ underlying capabilities around the tasks which require consistent temporal reasoning and deductions.
9 Limitations
The scope of TEMP-COFAC is to evaluate LLMs for their temporally consistent factuality capabilities. TEMP-COFAC comprehensively covers entity-entity temporal relations but falls short of stating entity-time and time-time aspects of temporal relations. Therefore, it should be applied along with other prominent datasets to estimate the overall temporal reasoning capabilities of LLMs.
On the other hand, the challenges related to computational assets to stack LLMs continue. Due to asset limitations, we may not utilize more extensive or commercially accessible LLMs to comprehensively evaluate on TeCFaP and approve on the off chance that the preferences of CoTSeLF are also advantageous to those models. Moreover, the smooth variant of CoTSeLF looks promising from its conception point of view but it could not outperform its discrete variant’s peer. We leave an in-depth ablation on smooth variants for future research.
Furthermore, despite efforts to apply higher quality standards, TEMP-COFAC relies on human annotation and is therefore prone to annotation errors. Moreover, the base language of TEMP-COFAC is English; therefore, it falls short in measuring consistent temporal factuality for other languages, particularly low-resource ones. Extensions to multilingual setting or resource-poor languages are left to future research.
10 Ethics Statement
The TEMP-COFAC is based on the Wikipedia and open world wide web knowledge sources. Wikipedia articles are licensed under a Creative Commons Attribution-ShareAlike 4.0 International License***https://en.wikipedia.org/wiki/Wikipedia:FAQ/Copyright (CC BY-SA 4.0) and its knowledge base is in the public domain. We will release TEMP-COFAC under same licence too. The experiments are conducted with all open source LLMs. Authors do not intend to introduce biases in any form to LLMs While applying fine-tuning methods.
The subjects and entities selected to be part of TEMP-COFAC are prone to unintended human biases during construction. The authors do not propagate any views/opinions, products, or representations of these subjects or entities in any form. The fact that women do not find representation during TEMP-COFAC annotations should be seen as a symptom of the gender disparity in research and innovation worldwide, but this is not the authors’ view. We support gender and racial equality in research and innovation with the utmost sincerity. Additionally, no generative AI-based content creation tools or applications were used to create this artifact, except for specialized support for spell checking, grammar correction, and paraphrasing.
Appendix A Appendix
A.1 Extended Description for TEMP-COFAC
In this section we continue to present details on TEMP-COFAC dataset. TEMP-COFAC is the first dataset of its kind in the temporal consistency domain, providing a strictly homogeneous sequence of entities for diverse subject-relation pairs. The conditioning on events that occurred without recurrence for strict temporal relationships along with the prefix formulation of key-object, sub-rel, value-object posed a severe challenge and required significant human evaluations at the level of subject-relation pairs during construction.
Figure 9: Entity types distribution in TEMP-COFAC resource as % of overall entities in data and representation across subject-relation (subject-relation) pairs.
Entity Types Distribution. We present the overall entity types distribution across different subject-relation pairs in Figure 9. TEMP-COFAC has broader representations for various entity types such as person names, countries, books, movie and album names, satellites, vehicles, and software.
Temporal Distribution. The temporal distribution of entities is presented in Table 4. We observe that the entities have a more comprehensive range of temporal representations in TEMP-COFAC dataset spanning in the range of 1500 to 2022. Most of the pre-training datasets that LLMs are trained on have a cutoff year of 2020, with few exceptions. Therefore, we consider that $99\%$ of all entities in TEMP-COFAC dataset must belong to a year equal to or less than 2020. It can be observed that there is a skewed temporal distribution of entities in favor of entities in the contemporary period (the year 1800-) compared to the historical period (the year 1500-1800). It is noted that $57\%$ of all entities are from 1991 to 2020.
Bins (year) #Entities Bins (year) #Entities
1521-1530 0.08 1881-1890 1.42
1531-1540 0.16 1891-1900 1.34
1551-1560 0.16 1901-1910 2.52
1601-1610 0.16 1911-1920 3.79
1621-1630 0.16 1921-1930 3.31
1651-1660 0.16 1931-1940 2.76
1701-1710 0.32 1941-1950 0.55
1711-1720 0.08 1951-1960 3.00
1731-1740 0.24 1961-1970 6.07
1741-1750 0.16 1971-1980 5.60
1751-1760 0.32 1981-1990 6.86
1761-1770 0.32 1991-2000 11.75
1781-1790 2.05 2001-2010 19.79
1791-1800 0.39 2011-2020 25.39
1871-1880 0.87 2021-2030 0.24
Table 4: Temporal distribution of entities (in %) in TEMP-COFAC with a bin size of $10$ years.
Evidences from Pre-training Data. TeCFaP is a novel task of evaluating the consistent temporal relationship between entities in LLMs. Here, we present a few manually extracted evidence from pre-training data to support that the objective of TeCFaP is fairly expected from LLMs. We consider Wikipedia for this test as it is a part of the pre-training dataset for most of the recent LLMs, including LLaMA. Evidences are manually extracted from Wikipedia for an entity pair hybrid-theory and meteora for a subject-relation pair Linkin Park and Release. Figure 10 presents the evidence with their sources in the pre-training dataset. Given these sentences, a human can easily find the temporal relation between hybrid-theory and meteora. Therefore, it is a fair ask from LLMs to learn the temporal relation between entities given such sentences.
Figure 10: A few extracts from Wikipedia for a subject-relation pair (linkin park - release-by). We find the presence of sentences in pre-training data (Wikipedia) of LLMs, which defines the temporal relationship among entities associated with the given subject-relation pair.
Annotations Quality. We ensure the high quality of TEMP-COFAC resource by applying it to execute cross-annotator agreement experiment. To assess the qualitative measure of factuality, We randomly select a hundred samples of filled patterns across different subject-relation pairs. The two reviewers split the data in half and reviewed the samples that the other annotator produced during construction. They score the agreements on a scale of 5-point (1-lowest agreement and 5-complete agreement on factuality of filled pattern) Likert scale Likert (1932). Similarly, to give the qualitative measure for consistency of patterns created, we again randomly sample hundreds of filled patterns and pair these positively with another paraphrased filled pattern sampled randomly from respective subject-relation. The reviewers repeat the similar scoring methodology in factuality assessment (1-lowest agreement and 5-complete agreement if the two patterns are paraphrased). We observe a mean agreement of $4.84\pm 0.39$ out of $5$ maximum for factuality and a mean agreement of $4.86\pm 0.49$ on similar lines for consistency.
TEMP-COFAC Coverage. We present the TEMP-COFAC comparison with prior datasets in Table 5. Some of the columns data reused from Tan et al. (2023) comparison of datasets. TEMPREASON is one of the most comprehensive temporal reasoning datasets, with 21K queries of event-event probe type in the QA setting. Here are the six types of entities it has considered during automatic construction from Wikipedia: Person, School, Political party, Company, Position, and Sports team. A significant drawback of TEMP-REASON is that all the queries are anchored around just person names, either as a subject or as an object. I.e., Which team did <subject> play for before/after oj? or Who was the head of the government of <subject> before/after oj?.
In comparison, the proposed TEMP-COFAC has 10.1K prefix-style queries and covers eleven diverse entity types from several domains: Movies, Geographical location, Games, Albums, Persons, Satellites, Software, Books, Vehicles, Songs, and Elements. Here, the queries are anchored around eleven entity types via sixty-six diverse subject-relation pairs, such as Linkin Park released Hybrid Theory just before ___ or The FIFA U-17 World Cup was hosted by Canada immediately after ___.
The dataset will facilitate the further development around various aspects of consistent temporal reasoning such as consistent event sequencing, multi-hop sequence-based QA, consistent multi-reasoning QA, and architectural study of LLMs through temporal explainability of inconsistent behaviour via paraphrased queries.
Dataset Format Knowledge Paraphrases Temporal Sequence of Time Size
Source Homogeneously Grouped Coverage
Entities
TEMPCOFAC Prefix-Style:Language Human Annotated/Open Web 1526-2022 10.1K
TEMPREASONTan et al. (2023) QA:Language Wikipedia/Wikipedia 634-2023 52.8K
TEMPLAMADhingra et al. (2022) QA:Language Wikipedia 2010-2020 50k
Time-SensitiveQAChen et al. (2021) QA:Language Wikipedia/Wikipedia 1367-2018 41.2k
StreamingQALiška et al. (2022) QA:Language WMT 2007-2020 147k
TempQuestionsJia et al. (2018b) QA:KG Freebase NA 1.2k
TimeQuestionsJia et al. (2021) QA:KG Wikipedia NA 16.1k
CronQuestionsSaxena et al. (2021) QA:KG Wikipedia 34-2021 410k
Table 5: Comparison of TEMPCOFAC with prior datasets.
A.2 Extended Results
A.2.1 ICL Setting Cont.
Here, we continue from the result section and present results for temporal-factuality and temporal-consistency in ICT setting (Figure 11).
(a) Temporal factuality
(b) Temporal consistency
Figure 11: Results for k-shot (k=1,2,3) ICL setup with LLaMA[13B] in an open vocabulary setup across – (a) Temp-fact: temporal factuality, (b) Temp-cons: temporal consistency.
A.2.2 Closed Vocabulary Setting Cont.
(a) Temporal Factuality
(b) Temporal Consistency
Figure 12: A comparison between open and closed (using restricted candidate set) vocabulary settings across variants of LLaMA model in zero-shot and in-context learning setup with k-shot (k=3).
We continue from the Section 5 to present the results for metrics temporal-factuality and temporal-consistency in closed vocabulary setting in Figure 12. Three-shot with closed vocab setting significantly outperforms by attaining maximum scores of $12.79\%$ and $55.06\%$ for temporal-factuality and temporal-consistency respectively across various variants and settings.
A.2.3 Error Analysis Cont.
The positive case is presented in Table 6 where CoTSeLF improves the temporal consistent factuality over a baseline model. It can be observed that The temporal consistency-driven factual improvement makes substantial changes both in the model temporal consistency and factuality.
S.No. Paraphrases Baseline Model $\mathbf{\texttt{CTSRL}_{Discrete}}$
1 the academy award for best picture was won by the departed immediately before no country for old men no country for old men
2 the departed won the best picture academy award just before the reader no country for old men
3 the departed won the oscar for best picture right before the curious case of benjamin no country for old men
4 the departed won the best picture academy award succeeded by the curious case of benjamin no country for old men
5 movie the departed won the academy award for best picture right before the curious case of benjamin the artist
6 the departed won the oscar for best picture immediately prior to the curious case of benjamin no country for old men
7 the academy award for best picture was won by the departed soon before the curious case of benjamin no country for old men
8 the departed was awarded the academy award for best picture immediately prior to the curious case of benjamin the departed
Table 6: A comparison of output generated by a baseline model and the discrete variant of CTSRL model for the sentence completion task. Where, subject-relation pair is "academy award for best picture - win by", key_object is "the departed," and an expected value_object is "no country for old men." The rest of the settings are similar as in the case of Table 3.
A.2.4 Probabilistic Space Analysis
We conduct an exploration for the evidences of improvement in model consistency through its probabilistic space analysis. The divergence in probability distribution for the next word shall be minimal for identical intent paraphrases to generate consistent value_object for given key_object. Similarly, the divergence shall be wider for different intent paraphrases. We use positive and agnostic paraphrases as notations to denote paraphrases with identical intent and paraphrases with different intent. We apply random hard negative sampling while considering the agnostic paraphrases. Here is an example of a positive and agnostic paraphrases.
Positive paraphrases
P1: hybrid theory was released by linkin park just before
P2: linkin park released hybrid theory immediately before
Agnostic paraphrases
P1: hybrid theory was released by linkin park just before
P2: linkin park released hybrid theory immediately after
The KL divergence metric is widely used to compare probabilistic distributions. We calculate the KL divergence of subsequent word’s probability distribution between positive and agnostic paraphrases, respectively. The objective is to maximize the difference between these two scores. We perform this experiment on all the entities for randomly selected ten subject-relation from test data. We randomly sample five pairs of positive and respective agnostic sentences. The scores are then averaged over a subject-relation pair, presented in Table 7. We compare the scores between the default LLaMA[13B] model and $\texttt{CTSRL}_{Discrete}$ model.
It is evident from Table 7 that the CoTSeLF improves the average difference between the KL divergence scores of positive paraphrase and agnostic paraphrase by the value of $0.18$ nats. We also observe a positive change in eight subject-relation pairs out of a total ten in the experiment. This analysis helps explain why the CoTSeLF achieves better temporally-consistent-factuality.
Subject-Relation Default $\mathbf{\texttt{CTSRL}_{Discrete}}$ $\Delta$
(ID) PP AP Diff (A) PP AP Diff (B) (A - B)
34 1.42 1.52 0.09 1.67 2.10 0.43 0.34
1 5.01 4.95 -0.06 5.51 5.19 -0.32 -0.26
19 1.99 2.01 0.03 2.77 2.91 0.14 0.11
42 1.79 1.72 -0.07 2.89 3.36 0.47 0.54
28 3.98 4.07 0.09 4.07 4.6 0.54 0.45
49 1.62 1.64 0.02 2.07 2.13 0.05 0.03
36 1.14 1.29 0.15 2.96 2.91 -0.05 -0.20
53 3.09 2.92 -0.17 2.26 2.17 -0.08 0.08
9 2.21 2.22 0.02 3.22 3.42 0.20 0.18
8 2.43 2.41 -0.02 2.15 2.65 0.50 0.51
Average 2.47 2.48 0.01 2.96 3.15 0.19 0.18
Table 7: The results present a comparison between the default LLaMA[13B] model and $\texttt{CTSRL}_{Discrete}$ variant (a fine-tuned LLaMA[13B] model) of CoTSeLF in probabilistic space. Where PP and AP are the KL divergence score between the next word’s probability distribution of positive paraphrases and agnostic paraphrases, respectively, Diff is the score difference between agnostic paraphrases (AP) and positive paraphrases (PP), and $\Delta$ is the difference between the two models differentials outcome. A positive value of $\Delta$ represents that the CoTSeLF has a broader separation of divergence in agnostic paraphrases compared to positive paraphrases, reflecting a more consistent model.
A.2.5 Significance Test for Alpha
We carry out the experiment on significance of parameter $\alpha$ used in $\texttt{CTSRL}_{Discrete}$ formulation. The experimental and data setting is the same as in section 5 for MT-IT + $\texttt{CTSRL}_{Discrete}$ formulation. The results are presented in Figure 13 for three different values of $\alpha$ (= 0.5, 0.66 and 0.75) across all three metrics temporal-factuality, temporal-consistency and temporally-consitent-factuality respectively. We observe the optimal performance at $\alpha$=0.66 for temporal-factuality and temporally-consitent-factuality. Whereas temporal-consistency further improves as we increase the value of $\alpha$. We have selected $\alpha$ as 0.66 for the main results presented in Table 3 based on the outcome of this ablation.
(a) Temporal Factuality
(b) Temporal Consistency
(c) Temporally Consistent Factuality
Figure 13: The results present the significance of a parameter alpha ($\alpha$) in the formulation of $\texttt{CTSRL}_{Discrete}$ across three metrics: (a) temporal factuality, (b) temporal consistency, and (c) temporally consistent factuality respectively in percentages. The rest of the settings for this experiment are the same as in the case of Table 3.
A.2.6 GPT-4 Evaluation Results
This section presents the evaluation results for a prominent commercial LLM, GPT-4. We conduct this experiment with a limited randomly selected 100 samples from the TEMP-COFAC due to financial limitations. We set up the probe in zero-shot, where the task is to complete a sentence with the correct phrase. To this purpose, an instruction (Instruction: "complete the given sentence with the correct phrase") is provided along with the input to the model. In a zero-shot setting, the temporal-factuality stands at $31.00\%$. Trained with a trillion of parameters, in absolute terms, the temporal-factuality in GPT-4 is not very impressive, although it surpasses the relatively mid-sized LLMs such as LLaMA with a substantial margin in the zero-shot setting. An example of each of the positive and negative responses is mentioned in Table 8. CoTSeLF’s significant improvements over LLaMA[13B] do not exceed the standard performance of GPT-4. This is expected given the large differences in their world knowledge through the respective parameter sizes and an access to the open internet in case of GPT-4. Furthermore, we could not assess CoTSeLF’s efficacy on GPT-4 model due to the inaccessibility of its architecture for open fine-tuning.
Due to a minimal sample set, we are short of reporting either the temporally-consistent-factuality or temporal-consistency for this test. Metrics like temporal consistency and temporally consistent factuality require a sufficient number of pair of samples for a subject-relation to calculate those. The other factors are temporal direction and entity pairs, which are to be considered while calculating. The 100 individual samples might be enough to gauge the preliminary assessment of temporal factuality but statistically insufficient for reporting temporally consistent factuality. We leave to improve the temporally-consistent-factuality of GPT-4 for future research via methods that may not require gradient updates. Due to a minimal sample set, we are short of reporting either the temporally-consistent-factuality or temporal-consistency for this test. Metrics like temporal consistency and temporally consistent factuality require a sufficient number of pair of samples for a subject-relation to calculate those. The other factors are temporal direction and entity pairs, which are to be considered while calculating. The 100 individual samples might be enough to gauge the preliminary assessment of temporal factuality but statistically insufficient for reporting temporally consistent factuality. We leave to improve the temporally-consistent-factuality of GPT-4 for future research via methods that may not require gradient updates.
S.No. Prompt Correct Answer Model’s Response
1 Black panther won the best costume design academy award succeeded by little women little women
2 The FIFA u-17 world cup was hosted by canada immediately after China Scotland
Table 8: An example of each of the positive and negative response from GPT-4 in zero-shot setting.
A.3 Hyperparameters
All codes were composed utilizing PyTorch. We utilized the Huggingface***https://huggingface.co/ repository for stacking the open-source LLMs. The implementation of RL is carried out through trlX, a python-based library Castricato et al. (2023). A PEFT-based method LoRA with 8-bit quantization is used for both instruction-tuned and PPO-based RL models. We apply minimal pre-processing on generated outputs from LLMs  such as filtering of articles or special characters (at max up to one word) before retrieving the value_object. The maximum sequence length of $256$ is used across all experiments. An NVIDIA A100 GPU (80 GB)***https://www.nvidia.com/en-in/data-center/a100/ was used to train the model.